OpenStudy (anonymous):

Which of the following is equivalent to the proportion 2/3=3/y? A. 3y = (2)(3) B. 2y = (2)(3) C. 2y = (3)(3) D. 3y = (3)(3)

OpenStudy (anonymous):

@jim_thompson5910 i stuck

OpenStudy (anonymous):

Hint: You can cross multiply here.

jimthompson5910 (jim_thompson5910):

when you cross multiply 2/3=3/y what do you get?

OpenStudy (anonymous):

no solution

jimthompson5910 (jim_thompson5910):

a/b = c/d cross multiplies to a*d = b*c

OpenStudy (anonymous):

y =9/2

jimthompson5910 (jim_thompson5910):

dont' completely solve for y though

jimthompson5910 (jim_thompson5910):

just do this first step of cross multiplication

OpenStudy (anonymous):

2y=9

jimthompson5910 (jim_thompson5910):

better, but that 9 can be written as what

OpenStudy (anonymous):

9x

jimthompson5910 (jim_thompson5910):

9 = 9*1 9 = something * something what goes in the place of "something" ?

OpenStudy (anonymous):

2y= (3)(3)

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Which of the following is equivalent to the proportion 3/4=x/12

jimthompson5910 (jim_thompson5910):

I'm assuming there are choices post them please and tell me which one you got

OpenStudy (anonymous):

A. 4/3= 12/x B. 3/4= 12/x C. 4/3= x/12 D. 4/12= x/3

jimthompson5910 (jim_thompson5910):

which one did you get

jimthompson5910 (jim_thompson5910):

hint: if a/b = c/d then you can flip both sides to get the equivalent equation b/a = d/c

OpenStudy (anonymous):

A right?

jimthompson5910 (jim_thompson5910):

good, you got it

OpenStudy (anonymous):

What is the first step that you have to do in order to solve the proportion: 1/3=y/26 A. 8.5 B. 26y = 3 C. y = 3(26) D. 3(y) = 1(26) I got B!:)

jimthompson5910 (jim_thompson5910):

B is incorrect

jimthompson5910 (jim_thompson5910):

a/b = c/d cross multiplies to a*d = b*c

OpenStudy (anonymous):

1/3=y/26 26=3y

jimthompson5910 (jim_thompson5910):

good, you got it, so that's close to 3(y) = 1(26) you just have to factor and flip the equation

OpenStudy (anonymous):

so our answer would be D then

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

When movies are being created, the set designers often use miniature models before they create the actual movie set. The set designers for Jurassic Park 3, created a 6 inch high and 4 inch long model dinosaur that represented a 20 feet high dinosaur on the set. How long was the actual dinosaur, rounded to the nearest foot? A. 15 feet B. 13 feet C. 14 feet D. 12 feet

OpenStudy (anonymous):

I think its B

jimthompson5910 (jim_thompson5910):

solve for x 6/20 = 4/x

OpenStudy (anonymous):

x=40/3

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

which is x = 13.333 in decimal form round this to the nearest whole number to get x = 13

OpenStudy (anonymous):

so our answer is 13

jimthompson5910 (jim_thompson5910):

you got it